TechNo stock was $25 per share at the end of last year since then it paid a $50 per share dividend last year to stock price is currently $23 if you won't 300 shares of TechNo what was your percent return
Answer #1

Percentage return = (Dividend + stock appreciation) / Stock price at beginning of year

Percentage return = {50 + (23-25)} / 25 = 48/25 = 192%